OpenStudy (anonymous):

cos 3A=4 cos^3 A-3 cos A how to prove it?

OpenStudy (pratyush5):

cos (3A) = cos(2A + A) = cos (2A) cos (A) - sin(2A) sin(A) = [ 2cos^2(A) - 1 ] cos (A) - (2 sin A cos A )sin A = 2cos^3(A) - cos A - 2sin^2(A) cos A = 2cos^3(A) - cos A - 2( 1 - cos^2(A)) cos A = 2cos^3(A) - cos A - 2cos A + 2cos^3(A) = 4cos^3(A) - 3cos A

OpenStudy (pratyush5):

You can also start like this cos (3A) = cos(A + 2A)

OpenStudy (pratyush5):

I believe you know this identity 

OpenStudy (anonymous):

yeah i know.. anyway thanks. u r genius man